# Where did heavy binaries go? Gravitational-wave populations using Delaunay triangulation with optimized complexity

### v2

@@ Abstract
We investigate the joint mass-redshift evolution of the binary black-hole merger rate in the latest gravitational-wave detection catalog, GWTC-4.0. We present and apply a novel non-parametric framework for modeling multi-dimensional, correlated distributions based on Delaunay triangulation. Crucially, the complexity of the model—namely, the number, positions, and weights of triangulation nodes—is inferred directly from the data, resulting in a highly efficient approach that requires about one to two orders of magnitude fewer parameters and significantly less calibration than current state-of-the-art methods. We find no evidence for a peak at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ at low redshifts ($z\sim 0.2$), where it would correspond to the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature reported in redshift-independent mass spectrum analyses, and we infer an increased merger rate at high redshifts ($z\sim 1$) around those masses, compatible with such a peak. When related to the time-delay distribution from progenitor formation to binary black-hole merger, our results suggest a short-delay, dense‑environment origin for sources contributing to the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature at high redshifts, and longer delays at lower redshifts compatible with isolated binary evolution.
software: astropy (Robitaille et al., 2013), eryn (Karnesis et al., 2023), numpy (Harris et al., 2020), qhull (Barber et al., 1996), scipy (Virtanen et al., 2020).
@@ 1 Introduction
The increasing number of binary black holes (BBHs) observed by gravitational-wave (GW) interferometers (Abbott et al., 2019, 2021, 2023a, 2024; Abac et al., 2025d) deepens our understanding of their population in the Universe (Abac et al., 2025c), eventually revealing their origin out of the multiple proposed formation channels (Mandel & Farmer, 2022; Mapelli, 2021). A crucial step in this direction is the identification of correlations between parameters and the presence of distinct subpopulations (e.g. Callister 2024 and references therein)
@@
A key question is whether, and how, the BBH population evolves with redshift and, crucially, whether this evolution depends on the BBH properties. BBHs are the end products of stellar evolution; therefore, their population should depend on the star-formation history and properties of galaxies such as metallicity (Mapelli et al., 2019; Neijssel et al., 2019; van Son et al., 2022; de Sá et al., 2024; Broekgaarden et al., 2021). Moreover, BBHs involving remnants of previous mergers, formed through dynamical encounters, also contribute to the evolution of the population’s properties (Gerosa & Fishbach, 2021; Ye & Fishbach, 2024; Torniamenti et al., 2024). Finally, the redshift evolution encodes the relative contribution of the different BBH formation channels throughout cosmic history (Zevin et al., 2021; Mapelli et al., 2022; Sedda et al., 2023).
@@
Evidence for an evolution of the effective spin distribution with redshift (Biscoveanu et al., 2022) was found on the third gravitational-wave transient catalog (GWTC-3) by LIGO, Virgo, and KAGRA, and was further strengthened in their fourth catalog (GWTC-4.0; Abac et al. 2025c). In contrast, several analyses of GWTC-3 found no evidence for or against a redshift evolution in the mass distribution (Ray et al., 2023; Heinzel et al., 2025b; Sadiq et al., 2025; Lalleman et al., 2025; Gennari et al., 2025), and first analyses on GWTC-4.0 find similar conclusions (Abac et al., 2025c). [^1] While Rinaldi et al. (2024) did report evidence for such an evolution, their treatment of selection effects is not accurate (Essick & Fishbach, 2024; Toubiana et al., 2025).
@@
In this work, we investigate the joint mass-redshift evolution of the BBH merger rate using a novel non-parametric multi-dimensional approach. In particular, we reconstruct the joint mass–redshift BBH merger rate using Delaunay triangulation (Delaunay, 1934) and barycentric interpolation. Both the number and locations of the triangulation vertices, as well as their associated weights, are inferred directly from the data through trans-dimensional Bayesian inference (Toubiana et al., 2023). The use of a data-driven interpolation scheme assumes no specific functional (in)dependence between mass and redshift; this allows us to probe a broader parameter space compared to models with closed-form correlations.
@@
Using GWTC-4.0 data, we report a distinct difference in the distribution of masses at $z=0.2$ and $z=1.0$ at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$, with no evidence for the presence of a peak in the merger rate at said masses for $z=0.2$. We then discuss the astrophysical implications of our findings.
@@ 2 Mass-redshift correlation
where $\Delta(M_{\mathrm{tot}},z|\Lambda_{\Delta})$ represents the (natural log) differential rate as computed by barycentric interpolation using Delaunay triangulation. The hyperparameters $\Lambda_{\Delta}$ are the number of triangulation vertices, their locations, and the value of the differential rate at those vertices. Figure 1 gives a schematic description of our model; further details are provided in Appendix A.
We model the mass ratio distribution as a broken power-law, which is compatible with the marginalised $q$ distribution found in Abac et al. (2025c, see Appendix B for details) For the spin magnitudes and tilt angles we assume the same functional forms as in the default model by Abac et al. (2025c). The hyperparameters $\Lambda=(\Lambda_{\Delta},\Lambda_{q},\Lambda_{\chi},\Lambda_{\vartheta})$ are inferred through hierarchical Bayesian inference (see Appendix C for details). The differential volumetric rate of GW events is then given by
@@
We limit the domain of inference for the $(M_{\mathrm{tot}},z)$ distribution to $6,350]\,\mathrm{M}_{\odot}\times[0,2.5]$. Priors on triangulation vertices are uniform within this domain, and weights are uniformly distributed along (20,15)(-20,15). The number of triangulation nodes is allowed to vary between 4 and 100. Inference is conducted using reversible-jump Markov chain Monte Carlo as implemented in eryn (Karnesis et al., 2023). We have verified our findings to be robust against the choice of prior (tests are reported in Appendices B and C).
@@
The right panel of Fig. 2 shows the credibility for the $z\sim 1$, $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ feature, which we quantify as the posterior probability of $\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z=1)>\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z=0.2)$. This value is $\gtrsim 95\%$ for $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$, and reaches a maximum of $98\%$ at $M_{\mathrm{tot}}\sim 63\,\mathrm{M}_{\odot}$. Said increase with redshift is compatible with that found by Abac et al. (2025d), where the merger rate is modeled as a function of redshift (but not mass) with an ansatz $\mathrm{d}_{\theta}\mathcal{R}\propto(1+z)^{\kappa}$ and returned $\kappa\sim 3$.
While we cannot definitively claim that the mass distribution at $z=1.0$ exhibits a peak at $\sim 70\,\mathrm{M}_{\odot}$ owing to the error bars, a corresponding feature at $m_{1}\sim 35\,\mathrm{M}_{\odot}$ has been consistently identified since GWTC-2 (Abbott et al., 2021). Our findings suggest that this feature is associated with high-redshift BBHs and disappears by $z=0.2$. In Appendix D, we show the results of applying our method to $(m_{1},z)$ and illustrate how our discussion of the $70\,\mathrm{M}_{\odot}$ feature similarly applies to the $35\,\mathrm{M}_{\odot}$ peak in $m_{1}$. Altogether, these results indicate that this high-mass feature in the BBH mass distribution (Abbott et al., 2021, 2023b; Abac et al., 2025c) varies with redshift. These results are also compatible with Rinaldi et al. (2025), who reported tenative evidence for a steeper redshift evolution associated to the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ peak than that observed for the power-law component in the GWTC-3 catalog. Future analyses should consider targeting such behavior with more strongly modeled approaches.
@@
We characterize the high-redshift appearance of the $\sim 70\,\mathrm{M}_{\odot}$ feature by computing Kullback-Leibler divergence ($D_{\mathrm{KL}}$; Kullback & Leibler 1951) between the posterior and the prior on the rate induced by our prior on $\Lambda_{\Delta}$ For reference, $D_{\mathrm{KL}}\lesssim 0.1\,\mathrm{nat}$ between two independent prior draws. As show in the upper left panel of Fig. 2, we find significantly larger values $D_{\mathrm{KL}}\sim 2.5\,\mathrm{nat}$ around $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$, showing that these results are primarily driven by the observed data. High $D_{\mathrm{KL}}$ values align with low relative uncertainties in Fig. 3 ($\mathrm{d}_{\theta}\mathcal{R}^{95\%}/\mathrm{d}_{\theta}\mathcal{R}^{5\%}\lesssim 10$), where the inference is dominated by the observed events.
@@
Our results differ from those by Abac et al. (2025c), which reported no evidence for a mass-dependent redshift evolution of the merger rate using both copulas (Adamcewicz & Thrane, 2022) and binned Gaussian processes (Ray et al., 2023). Copulas are suited to identify linear correlations, but not the kind of multi-modal structures inferred in Fig. 3. In the binned Gaussian process approach, the high-mass feature appears less pronounced at lower redshift, consistent with our findings, though large error bars blur the trend. Their results were shown as probability density functions of $m_{1}$ across redshift bins, which are normalized by the total number of events, a very uncertain quantity. As a result, error bars end up broadening. This issue is avoided when working with the volumetric merger rate. The counterpart limitation is that the nonoverlap of confidence bands of the volumetric differential rate at different redshifts is not, by itself, sufficient to establish redshift evolution of the mass distribution. This is shown in the right panel of Fig. 2, where a straight line remains consistent with the 90% credibility interval of the rate ratio, which is thus compatible with an overall, rigid rescaling of the $\mathrm{d}_{\theta}\mathcal{R}$. Finally, unlike the binned method, our interpolation points are data-driven rather than fixed, potentially capturing finer features by construction.
@@
We also infer a peak in the merger rate at $M_{\mathrm{tot}}\sim 20\,\mathrm{M}_{\odot}$ for $z=0.2$, but the lack of detections at higher redshifts prevents us from confirming or disproving its existence for $z\gtrsim 0.2$. At the high-mass end, the bump at $z=0.2$ around $M_{\mathrm{tot}}\sim 250\,\mathrm{M}_{\odot}$ is likely driven by a single event, GW231123 (Abac et al., 2025a), while the the apparent increase in the rate for $M_{\mathrm{tot}}\sim 150\,\mathrm{M}_{\odot}$ at $z=1$ aligns with the inferred source properties of GW190521 (Abbott et al., 2020). We caution against overinterpreting such features, which indeed correspond to low values of $D_{\rm KL}\sim 1\,{\rm nat}$.
@@
Our inference on the mass-ratio distribution is consistent with a larger density at $q\gtrsim 0.6$, similarly to what was reported by Abac et al. (2025c). Details are reported in Appendix B.
@@
The triangulation scheme that best fits the $(M_{\mathrm{tot}},z)$ portion of the population has ${17}_{-11}^{+25}$ nodes, corresponding to $56^{+75}_{-33}$ free parameters (90% credible interval). This is one to two orders of magnitude fewer than those needed by Heinzel et al. (2025b), and requires less tuning compared to Ray et al. (2023), which computes the final number of bins by re-running the analysis using increasingly finer grids. Crucially, both of those approaches rely on a fixed two-dimensional grid and are therefore expected to scale poorly with the number of dimensions, whereas Delaunay triangulation is naturally applicable to higher-dimensional correlations. The method proposed by Payne & Thrane (2023) and Guttman et al. (2025) requires a comparable number of parameters, but returns the maximum likelihood histogram of data without an associated error. The approach of Sadiq et al. (2022) seems to suffer from edge effects where data is scarce or absent, while our method correctly returns the prior in these regions of parameter space.
@@ 3 Astrophysical interpretation
Our triangulation-based reconstruction suggests that the known feature in the BBH mass spectrum at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$   ($m_{1}\sim 35\,\mathrm{M}_{\odot}$) is primarily due to mergers at $z\gtrsim 0.7$ and fades by $z\sim 0.2$. More specifically, as shown in Fig. 4, the evolution of $\mathrm{d}_{\theta}\mathcal{R}$ as a function of cosmic time $t$ between $z=1$ and $z=0.2$ appears to follow two regimes, with $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-1}$ at low redshift and a much steeper evolution closer to $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-4}$ at higher redshift. We stress these are indicative trends and not rigourous fits, which would require more data.
@@
At low redshift, $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-1}$ is compatible with typical predictions for isolated binary evolution, including sources ejected from stellar clusters  (see e.g. Dominik et al., 2012; Belczynski et al., 2016; Eldridge & Stanway, 2016; Lamberts et al., 2016; Mapelli et al., 2017).
By contrast, the steeper behavior at higher redshift $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-4}$ is in line with numerical simulations of cluster evolution where BBHs merge inside their host clusters (e.g. Rodriguez & Loeb, 2018; Sedda et al., 2024). Moreover, these dynamically assembled BBHs are, on average, more massive than those from isolated evolution  (Rodriguez et al., 2016; Di Carlo et al., 2019; Antonini & Gieles, 2020; Sedda et al., 2024), which would explain their prominence at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$   ($m_{1}\sim 35\,\mathrm{M}_{\odot}$(Antonini et al., 2023; Bruel et al., 2025).
@@
Massive, dense globular clusters match the requirements for this explanation, as they likely tend to form at earlier cosmic times with respect to the global formation of stars in the Universe (Choksi et al., 2018; El-Badry et al., 2019) and thus naturally give rise to a high-mass contribution to the BBH merger rate that switches off at low redshifts. Numerical analyses of this scenario predict a high-mass BBH peak at $z\gtrsim 2$ that extends to lower redshifts (Ye & Fishbach, 2024; Torniamenti et al., 2024), suggesting that our observation could in fact correspond to the lower-redshift tail of this peak. Similarly, AGN disks are another dense environment favoring hierarchical mergers which can also produce heavy BBHs with short delays (Yang et al., 2019; Tagawa et al., 2020; Santini et al., 2023).
@@
Pulsational pair-instability supernovae (PPISNe) have long been invoked to explain the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature as a pile-up of masses at the instability onset (see e.g. Woosley, 2017; Talbot & Thrane, 2018). Recent studies, however, have shown this hypothesis to be in conflict with theoretical predictions of the expected location of the related mass gap or the relative number of systems in this mass range (Stevenson et al., 2019; Farag et al., 2022; Hendriks et al., 2023). Moreover, the PPISN channel would require extreme values of nuclear reaction rates, in particular the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate, to match the BH mass distribution inferred from GW observations (see e.g. Tong et al., 2025; Antonini et al., 2025, for recent analyses). Also, while the interplay between time delays and the dependence of BH maximum mass on progenitor metallicity can give rise to an evolution with redshift of the BH mass spectrum (Farmer et al., 2019; Mukherjee, 2022), this scenario relies preferentially on longer time delays, which are not favored by our analysis. The results presented in this work add to the mounting evidence that this formation channel may not be the preferred one to explain the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature in the GW spectrum.
@@
Peaks in the BBH mass spectrum at $m_{1}\gtrsim 30\,\mathrm{M}_{\odot}$ can also be caused by low-metallicity stars whose structure has not been altered by binary interaction (Schneider et al., 2021, 2023), which then get paired through other processes. The low-metallicity requirement naturally ties this process to redshift through the chemical enrichment of the Universe, but further work is required to assess the compatibility of these models with our results.
@@
We also recover a low-mass peak $M_{\mathrm{tot}}\sim 20\,\mathrm{M}_{\odot}$ $(m_{1}\sim 10\,\mathrm{M}_{\odot})$ at $z=0.2$ (current uncertainties prevent confirmation at $z=1$). This previously reported feature (Abbott et al., 2021, 2023b; Abac et al., 2025c) is well-explained by isolated binary evolution at near-solar metallicity (Fragos et al., 2023; Agrawal et al., 2023) or dynamical evolution of stars and compact objects in metal-rich dense clusters (Ye et al., 2025). More observations are needed to confirm the invariance of this low-mass peak across redshift.

### v3

@@ Abstract
We investigate the joint mass-redshift evolution of the binary black-hole merger rate in the latest Gravitational-Wave Transient Catalog, GWTC-4.0. We present and apply a novel non-parametric framework for modeling multi-dimensional, correlated distributions based on Delaunay triangulation. Crucially, the complexity of the model—namely, the number, positions, and weights of triangulation nodes—is inferred directly from the data, resulting in a highly efficient approach that requires about one to two orders of magnitude fewer parameters and significantly less calibration than current state-of-the-art methods. We find no evidence for a peak at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ at low redshifts ($z\sim 0.2$), where it would correspond to the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature reported in redshift-independent mass spectrum analyses, and we infer an increased merger rate at high redshifts ($z\sim 1$) around those masses, compatible with such a peak. When related to the time-delay distribution from progenitor formation to binary black-hole merger, our results suggest that sources contributing to the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature follow a steeper (shallower) time-delay distribution at high (low) redshifts. This hints at contributions from different formation channels—for example dense environments and isolated binary evolution, respectively—although firm identification of specific formation pathways will require further observations and analyses.
software: astropy (Astropy:2013muo), eryn (Karnesis:2023ras), numpy (Harris:2020xlr), qhull (Barber:1996lmi), scipy (Virtanen:2019joe).
@@ 1 Introduction
The increasing number of binary black holes (BBHs) observed by gravitational-wave (GW) interferometers (LIGOScientific:2018mvr; LIGOScientific:2020ibl; KAGRA:2021vkt; LIGOScientific:2021usb; LIGOScientific:2025slb) deepens our understanding of their population in the Universe (LIGOScientific:2025pvj), eventually revealing their origin out of the multiple proposed formation channels (Mandel:2018hfr; Mapelli:2021taw). A crucial step in this direction is the identification of correlations between parameters and the presence of distinct subpopulations (e.g. Callister:2024cdx and references therein)
@@
A key question is whether, and how, the BBH population evolves with redshift and, crucially, whether this evolution depends on the BBH properties. BBHs are the end products of stellar evolution; therefore, their population should depend on the star-formation history and properties of galaxies such as metallicity (Mapelli:2019bnp; Neijssel:2019irh; vanSon:2021zpk; deSa:2024otm; Broekgaarden:2021iew). Moreover, BBHs involving remnants of previous mergers, formed through dynamical encounters, also contribute to the evolution of the population’s properties (Gerosa:2021mno; Ye:2024ypm; Torniamenti:2024uxl). Finally, the redshift evolution encodes the relative contribution of the different BBH formation channels throughout cosmic history (Zevin:2020gbd; Mapelli:2021gyv; Sedda:2021vjh).
@@
Evidence for an evolution of the effective spin distribution with redshift (Biscoveanu:2022qac) was found on the third gravitational-wave transient catalog (GWTC-3) by LIGO, Virgo, and KAGRA, and was further strengthened in their fourth catalog (GWTC-4.0; LIGOScientific:2025pvj). In contrast, several analyses of GWTC-3 found no evidence for or against a redshift evolution in the mass distribution (Ray:2023upk; Heinzel:2024hva; Sadiq:2025aog; Lalleman:2025xcs; Gennari:2025nho), and first analyses on GWTC-4.0 find similar conclusions (LIGOScientific:2025pvj). [^1] While Rinaldi:2023bbd did report evidence for such an evolution, their treatment of selection effects is not accurate (Essick:2023upv; Toubiana:2025syw).
@@
In this work, we investigate the joint mass-redshift evolution of the BBH merger rate using a novel non-parametric multi-dimensional approach. In particular, we reconstruct the joint mass–redshift BBH merger rate using Delaunay triangulation (Delaunay_1934aa) and barycentric interpolation. Both the number and locations of the triangulation vertices, as well as their associated weights, are inferred directly from the data through trans-dimensional Bayesian inference (Toubiana:2023egi). The use of a data-driven interpolation scheme assumes no specific functional (in)dependence between mass and redshift; this allows us to probe a broader parameter space compared to models with closed-form correlations.
@@
Using GWTC-4.0 data [^2] We analyze the 153 BBH events considered by LIGOScientific:2025pvj, see their Sec. 6. Note this includes only those events with less than 1% posterior support for component masses below $3\,\mathrm{M}_{\odot}$. , we report a distinct difference in the distribution of masses at $z=0.2$ and $z=1.0$ at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$, with no evidence for the presence of a peak in the merger rate at said masses for $z=0.2$. We then discuss the astrophysical implications of our findings.
@@ 2 Mass-redshift correlation
where $\Delta(M_{\mathrm{tot}},z|\Lambda_{\Delta})$ represents the (natural log) differential rate as computed by barycentric interpolation using Delaunay triangulation. The hyperparameters $\Lambda_{\Delta}$ are the number of triangulation vertices, their locations, and the value of the differential rate at those vertices. Figure 1 gives a schematic description of our model; further details are provided in Appendix A. We model the mass ratio distribution as a broken power-law, which is compatible with the marginalised $q$ distribution found in LIGOScientific:2025pvj. For the spin magnitudes and tilt angles we assume the same functional forms as in the default model by LIGOScientific:2025pvj. The hyperparameters $\Lambda=(\Lambda_{\Delta},\Lambda_{q},\Lambda_{\chi},\Lambda_{\vartheta})$ are inferred through hierarchical Bayesian inference (see Appendix C for details). The differential volumetric rate of GW events is then given by
@@
We limit the domain of inference for the $(M_{\mathrm{tot}},z)$ distribution to $6,350]\,\mathrm{M}_{\odot}\times[0,2.5]$. Priors on triangulation vertices are uniform within this domain, and weights are uniformly distributed along (20,15)(-20,15). The number of triangulation nodes is allowed to vary between 4 and 100. Inference is conducted using reversible-jump Markov chain Monte Carlo as implemented in eryn (Karnesis:2023ras). We have verified our findings to be robust against the choice of prior (tests are reported in Appendices B and C).
@@
The right panel of Fig. 2 shows the credibility for the $z\sim 1$, $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ feature, which we quantify as the posterior probability of $\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z=1)>\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z=0.2)$. This value is $\gtrsim 95\%$ for $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$, and reaches a maximum of $98\%$ at $M_{\mathrm{tot}}\sim 63\,\mathrm{M}_{\odot}$. Said increase with redshift is compatible with that found by LIGOScientific:2025slb, where the merger rate is modeled as a function of redshift (but not mass) with an ansatz $\mathrm{d}_{\theta}\mathcal{R}\propto(1+z)^{\kappa}$ and returned $\kappa\sim 3$.
While we cannot definitively claim that the mass distribution at $z=1.0$ exhibits a peak at $\sim 70\,\mathrm{M}_{\odot}$ owing to the error bars, a corresponding feature at $m_{1}\sim 35\,\mathrm{M}_{\odot}$ has been consistently identified since GWTC-2 (LIGOScientific:2020kqk). Our findings suggest that this feature is associated with high-redshift BBHs and disappears by $z=0.2$. In Appendix D, we show the results of applying our method to $(m_{1},z)$ and illustrate how our discussion of the $70\,\mathrm{M}_{\odot}$ feature similarly applies to the $35\,\mathrm{M}_{\odot}$ peak in $m_{1}$. Altogether, these results indicate that this high-mass feature in the BBH mass distribution (LIGOScientific:2020kqk; KAGRA:2021duu; LIGOScientific:2025pvj) varies with redshift. These results are also compatible with Rinaldi:2025emt, who reported tentative evidence for a steeper redshift evolution associated to the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ peak than that observed for the power-law component in the GWTC-3 catalog. Future analyses should consider targeting such behavior with more strongly modeled approaches.
@@
We characterize the high-redshift appearance of the $\sim 70\,\mathrm{M}_{\odot}$ feature by computing Kullback-Leibler divergence ($D_{\mathrm{KL}}$; kl_div) between the posterior and the prior on the rate induced by our prior on $\Lambda_{\Delta}$ For reference, $D_{\mathrm{KL}}\lesssim 0.1\,\mathrm{nat}$ between two independent prior draws. As show in the upper left panel of Fig. 2, we find significantly larger values $D_{\mathrm{KL}}\sim 2.5\,\mathrm{nat}$ around $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$, showing that these results are primarily driven by the observed data. High $D_{\mathrm{KL}}$ values align with low relative uncertainties in Fig. 3 ($\mathrm{d}_{\theta}\mathcal{R}^{95\%}/\mathrm{d}_{\theta}\mathcal{R}^{5\%}\lesssim 10$), where the inference is dominated by the observed events.
@@
Our results differ from those by LIGOScientific:2025pvj, which reported no evidence for a mass-dependent redshift evolution of the merger rate using both copulas (Adamcewicz:2022hce) and binned Gaussian processes (Ray:2023upk). Copulas are suited to identify linear correlations, but not the kind of multi-modal structures inferred in Fig. 3. In the binned Gaussian process approach, the high-mass feature appears less pronounced at lower redshift, consistent with our findings, though large error bars blur the trend. Their results were shown as probability density functions of $m_{1}$ across redshift bins, which are normalized by the total number of events, a very uncertain quantity. As a result, error bars end up broadening. This issue is avoided when working with the volumetric merger rate. The counterpart limitation is that the nonoverlap of confidence bands of the volumetric differential rate at different redshifts is not, by itself, sufficient to establish redshift evolution of the mass distribution. This is shown in the right panel of Fig. 2, where a straight line remains consistent with the 90% credibility interval of the rate ratio, which is thus compatible with an overall, rigid rescaling of the $\mathrm{d}_{\theta}\mathcal{R}$. Finally, unlike the binned method, our interpolation points are data-driven rather than fixed, potentially capturing finer features by construction.
@@
We also infer a peak in the merger rate at $M_{\mathrm{tot}}\sim 20\,\mathrm{M}_{\odot}$ for $z=0.2$, but the lack of detections at higher redshifts prevents us from confirming or disproving its existence for $z\gtrsim 0.2$. At the high-mass end, the bump at $z=0.2$ around $M_{\mathrm{tot}}\sim 250\,\mathrm{M}_{\odot}$ is likely driven by a single event, GW231123 (LIGOScientific:2025rsn), while the the apparent increase in the rate for $M_{\mathrm{tot}}\sim 150\,\mathrm{M}_{\odot}$ at $z=1$ aligns with the inferred source properties of GW190521 (LIGOScientific:2020iuh). We caution against overinterpreting such features, which indeed correspond to low values of $D_{\rm KL}\sim 1\,{\rm nat}$.
@@
Our inference on the mass-ratio distribution is consistent with a larger density at $q\gtrsim 0.6$, similarly to what was reported by LIGOScientific:2025pvj. Details are reported in Appendix B.
@@
The triangulation scheme that best fits the $(M_{\mathrm{tot}},z)$ portion of the population has ${17}_{-11}^{+25}$ nodes, corresponding to $56^{+75}_{-33}$ free parameters (90% credible interval). This is one to two orders of magnitude fewer than those needed by Heinzel:2024hva, and requires less tuning compared to Ray:2023upk, which computes the final number of bins by re-running the analysis using increasingly finer grids. Crucially, both of those approaches rely on a fixed two-dimensional grid and are therefore expected to scale poorly with the number of dimensions, whereas Delaunay triangulation is naturally applicable to higher-dimensional correlations. The method proposed by Payne:2022xan and Guttman:2025jkv requires a comparable number of parameters, but returns the maximum likelihood histogram of data without an associated error. The approach of Sadiq:2021fin, while having a very small number of free parameters, seems to suffer from edge effects where data is scarce or absent; our method correctly returns the prior in these regions of parameter space.
@@ 3 Astrophysical interpretation
Our triangulation-based reconstruction suggests that the known feature in the BBH mass spectrum at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$   ($m_{1}\sim 35\,\mathrm{M}_{\odot}$) is primarily due to mergers at $z\gtrsim 0.7$ and fades by $z\sim 0.2$. More specifically, as shown in Fig. 4, the evolution of $\mathrm{d}_{\theta}\mathcal{R}$ as a function of cosmic time $t$ between $z=1$ and $z=0.2$ appears to follow two regimes, with $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-1}$ at low redshift and a much steeper evolution closer to $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-4}$ at higher redshift. We stress these are indicative trends and not rigorous fits, which would require more data.
@@
At low redshift, $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-1}$ is compatible with typical predictions for isolated binary evolution, including sources ejected from stellar clusters  (see e.g. Dominik:2012kk; Belczynski:2016obo; Eldridge:2016ymr; Lamberts:2016txh; Mapelli:2017hqk).
By contrast, the steeper behavior at higher redshift $\mathrm{d}_{\theta}\mathcal{R}(t)\propto t^{-4}$ is in line with numerical simulations of cluster evolution where BBHs merge inside their host clusters (e.g. Rodriguez:2018rmd; Sedda:2023qlx). Moreover, these dynamically assembled BBHs are, on average, more massive than those from isolated evolution  (Rodriguez:2016kxx; DiCarlo:2019pmf; Antonini:2020xnd; Sedda:2023qlx), which would explain their prominence at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$   ($m_{1}\sim 35\,\mathrm{M}_{\odot}$(Antonini:2022vib; Bruel:2025sdq).
@@
Massive, dense globular clusters match the requirements for this explanation, as they likely tend to form at earlier cosmic times with respect to the global formation of stars in the Universe (2018MNRAS.480.2343C; 2019MNRAS.482.4528E) and thus naturally give rise to a high-mass contribution to the BBH merger rate that switches off at low redshifts. Numerical analyses of this scenario predict a high-mass BBH peak at $z\gtrsim 2$ that extends to lower redshifts (Ye:2024ypm; Torniamenti:2024uxl), suggesting that our observation could in fact correspond to the lower-redshift tail of this peak. Similarly, AGN disks are another dense environment favoring hierarchical mergers which can also produce heavy BBHs with short delays (Yang:2019cbr; Tagawa:2019osr; Santini:2023ukl).
@@
Pulsational pair-instability supernovae (PPISNe) have long been invoked to explain the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature as a pile-up of masses at the instability onset (see e.g. Woosley:2016hmi; Talbot:2018cva). Recent studies, however, have shown this hypothesis to be in conflict with theoretical predictions of the expected location of the related mass gap or the relative number of systems in this mass range (2019ApJ...882..121S; Farag:2022jcc; Hendriks:2023yrw). Moreover, the PPISN channel would require extreme values of nuclear reaction rates, in particular the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate, to match the BH mass distribution inferred from GW observations (see e.g. Tong:2025wpz; Antonini:2025ilj, for recent analyses). Also, while the interplay between time delays and the dependence of BH maximum mass on progenitor metallicity can give rise to an evolution with redshift of the BH mass spectrum (2019ApJ...887...53F; Mukherjee:2021rtw), this scenario relies preferentially on longer time delays, which are not favored by our analysis. The results presented in this work add to the mounting evidence that this formation channel may not be the preferred one to explain the $m_{1}\sim 35\,\mathrm{M}_{\odot}$ feature in the GW spectrum.
@@
Peaks in the BBH mass spectrum at $m_{1}\gtrsim 30\,\mathrm{M}_{\odot}$ can also be caused by low-metallicity stars whose structure has not been altered by binary interaction (Schneider:2020vvh; Schneider:2023mxe), which then get paired through other processes. The low-metallicity requirement naturally ties this process to redshift through the chemical enrichment of the Universe, but further work is required to assess the compatibility of these models with our results.
@@
We also recover a low-mass peak $M_{\mathrm{tot}}\sim 20\,\mathrm{M}_{\odot}$ $(m_{1}\sim 10\,\mathrm{M}_{\odot})$ at $z=0.2$ (current uncertainties prevent confirmation at $z=1$). This previously reported feature (LIGOScientific:2020kqk; KAGRA:2021duu; LIGOScientific:2025pvj) is well-explained by isolated binary evolution at near-solar metallicity (Fragos:2022jik; Agrawal:2023ecm) or dynamical evolution of stars and compact objects in metal-rich dense clusters (Ye:2025ano). More observations are needed to confirm the invariance of this low-mass peak across redshift.